OpenStudy (anonymous):

which linear equation represents a line containing the point (1,3)?

OpenStudy (anonymous):

do you have another point or the slope.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

y- 3 = m(x-1) , covers every line that goes through that point. m is the slope.

OpenStudy (anonymous):

well iknow that because of y=mx + b but that doesnt answer the question. theres actual answers

OpenStudy (anonymous):

do you mean its multiple choice?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

alright find the one where the slope is m and b = -m +3

OpenStudy (anonymous):

the answers dont have those numbers in it thought. example: one of the answers is x+2y=5

OpenStudy (anonymous):

1+ 2(3) = 7 it can't be that one. plug (x =1 and y =3 into the others and see which one makes sense.

OpenStudy (anonymous):

i got it ! thankyou